# ProbeMDE: Uncertainty-Guided Active Proprioception for Monocular Depth Estimation in Surgical Robotics

### v2

@@ I Introduction
Depth perception is a fundamental requirement for robotic systems operating in unstructured and deformable environments such as surgical settings. Monocular depth estimation (MDE) offers one potential solution due to its ability to infer dense 3D scene structure from a single RGB camera without the need for specialized depth sensors [ming2021deep]. However, MDE models often struggle with ambiguities arising from textureless surfaces, specular reflections, and occlusions, all of which are especially prevalent in robotic surgical settings [daher2025shades].
One avenue to address these limitations is to combine vision-based predictions with proprioceptive measurements [ferguson2020toward][Kumar_Hamlyn2025]. We define robotic proprioception as a robot’s ability to sense contact with external objects in addition to an awareness of its own position in space through tool-tip kinematics. Utilizing proprioception, we can provide sparse ground-truth depth measurements at contact points between the robot and the environment. While these measurements can provide accurate estimates of depth at specified locations, their acquisition has a cost as each proprioceptive measurement requires physical interaction with the environment that will slow down task execution. Thus, we seek to minimize the number of required measurements while maximizing the reconstruction accuracy of the scene.
@@ II-A Monocular Depth Estimation
Monocular depth estimation uses an image from a single camera viewpoint to predict the distance from the camera to each point in the 3D scene. State-of-the-art MDE methods produce depth maps with outstanding accuracy for a wide variety of everyday indoor and outdoor scenes [yang_depth_2024], [ke_repurposing_2024]. However, endoscopic videos of surgical scenes are out of distribution from the large-scale datasets these foundation models are trained on, resulting in suboptimal performance [shao_endomust_2025].
Efforts to improve MDE in endoscopy include adapting or fine-tuning foundation models with simulation data [shao_endomust_2025], [tan_accurate_2025], [li_monocular_2025], using a style transfer network to make simulation renderings more photo-realistic [jeong_depth_2024], or utilizing point cloud data from structure from motion as sparse-ground-truth depth cues [liu_dense_2020]. MDE models focused on a specific domain frequently use an adaptation of the U-Net architecture [ronneberger2015u], [jeong_depth_2024]. In this work, we draw from a similar toolbox of approaches while adding the ability to adaptively improve depth prediction through uncertainty-guided active sensing. As in prior works, we utilize a style-transfer network to make simulation RGB renderings more realistic. We then use this data to fine-tune a large foundation model, Depth Anything V2 [yang_depth_2024], which we subsequently use to generate pseudo-ground-truth data for training a model with better real-world performance. We introduce a novel source of cues for endoscopic MDE, robotic proprioception at specific locations within the scene which will maximally resolve the uncertainty of our prediction.
@@ II-B Uncertainty Quantification
As deep neural networks become more complex and less interpretable, uncertainty quantification has become a critical component of any such system, especially for safety-critical tasks such as robotic surgery. Many methods exist for quantifying predictive uncertainty in deep learning including deep ensembles [lee2015m], Monte Carlo dropout [milanes2021monte], [camarasa2020quantitative], [mae2021uncertainty], and test-time augmentation [wang_aleatoric_2019]. A comparison of these methods shows that deep ensembles consistently provide the most accurate quantification of uncertainty under distribution shift in the data [ovadia_can_2019]. An ensemble is created by training multiple models with different random weight initializations such that they each converge to a different local minima in the loss space. At inference time, we can use the variance of the ensemble members’ predictions as a proxy for uncertainty  [fort2019deep][pearce2018high][lakshminarayanan2017simple]. Many works have quantified the uncertainty of MDE predictions [xiang_measuring_2024][marsal_monoprob_2024][landgraf_critical_2025], including some focused specifically on endoscopic images [chong_monocular_2025]. However, no past works have explored the active refinement of MDE predictions using robotic touch sensing. In this work, we leverage uncertainty quantification techniques for ensembles to develop an uncertainty-guided active touch sensing framework for monocular depth estimation.
@@ II-C Active Touch Sensing
Active touch sensing has been widely explored as a means of acquiring task-relevant information under sensing constraints. Early approaches focused on reconstructing object geometry through sparse and targeted contact interactions For instance, tactile fingertips have been used to actively trace object contours for shape recognition [assaf2014seeing], while force-based probing has enabled reconstruction of unknown objects [kim2015exploration]. More recent work has emphasized uncertainty-driven exploration, where sensing actions are guided to reduce ambiguity [martinez2017active]. Recent research has further advanced active touch sensing with machine learning and task-specific exploration. Learning-based approaches have been proposed to determine not only how to explore unknown objects, but also learning where to touch and how to recognize objects from tactile data [shahidzadeh2024actexplore, xu2022tandem]. Together, these works demonstrate the value of active touch sensing to improve perception and task performance. In contrast, our work extends the concept of active touch sensing into the domain of proprioceptive depth acquisition, where sparse depth measurements are strategically chosen to reduce uncertainty in monocular depth estimation for deformable surgical environments.
@@ IV-A Monocular Depth Estimation
Our MDE is based on a U-Net architecture which takes as input the RGB image of a surgical scene as well as a sparse set of ground-truth depth measurements [ronneberger2015u]. The sparse measurements are represented as a depth map where unobserved pixels are set to -1, enabling the network to condition its predictions on available propriocepted depths. The MDE then predicts a full depth map of the scene, leveraging both visual features in the RGB image as well as the sparse depth information.
@@ IV-C Active Depth Sensing
After computing the gradient of the total ensemble uncertainty with respect to the sparse depth measurements, $\nabla_{\mathrm{S}}U_{\mathrm{total}}$, we use this information to actively select which locations to propriocept. Rather than selecting points greedily, we adopt Stein Variational Gradient Descent (SVGD) to optimize a set of candidate locations [liu2016stein]. Selecting points greedily could lead to a set of query points all near each other. SVGD strikes a balance between clustering points in highly informative regions while also encouraging diversity throughout the image. We first negate the gradient map so that the maximum values in the map correspond with locations with the highest likelihood of reducing uncertainty. The resulting map is then normalized to form a discrete probability distribution over all candidate pixel locations, indicating where measurements would be the most informative.
@@ V-A Datasets
Simulated Dataset: A simulated dataset is generated from CT scans of 13 real-world ex vivo porcine tracheal models with an intraluminal tumor surrogate (chicken tissue) affixed using cyanoacrylate adhesive. We convert the scans into 3D meshes in Unity. A simulated camera explores the interior of the trachea reconstructions, collecting pairs of RGB images and depth maps. The RGB images are then translated into the style of real phantom endoscopic scenes using a style-transfer network [park2020contrastive] to reduce the domain gap.
Real-world Dataset: Our real-world dataset is collected from 12 other ex vivo porcine tracheal models with tumor surrogates prepared as above. We collect exploratory endoscopic videos of these models using the Virtuoso Endoscopy System (Virtuoso, Nashville, TN, USA)  [virtuoso-ves-manual] and then generate pseudo-ground-truth depth maps using Depth Anything V2’s indoor large model [yang_depth_2024] fine-tuned on the simulated dataset. We create additional data points by linearly and exponentially stretching the values of the pseudo-ground-truth depth maps. This artificially exaggerates the ambiguity of perceived scale, thereby forcing the model to rely heavily on the sparse ground truth.
@@ V-C Experimental Setup
Whereas in the simulated setting, a dense ground-truth depth map was readily available, this information does not exist in the physical system. Consequently, depth estimates are obtained through a touch-sensing tool integrated at the distal end of the concentric tube arm. The touch-sensing tool, first introduced in [Kumar_Hamlyn2025], employs a capacitance circuit coupled to an electrosurgery probe running along the arm. Tissue contact produces a sharp change in capacitance, which can be measured in real time, thereby enabling the arm to detect contact events. By using this tool to physically ‘probe’ image-space points, we are able to recover their corresponding depths via robot kinematics.
@@ V-E Evaluation Metrics
To evaluate the reconstruction quality of each method, we report standard monocular depth estimation metrics widely used in prior work [eigen2014depth, eigen2015predicting, godard2017unsupervised, laina2016deeper]:

### v3

@@ I Introduction
Depth perception is a fundamental requirement for robotic systems operating in unstructured and deformable environments such as surgical settings. Monocular depth estimation (MDE) offers one potential solution due to its ability to infer dense 3D scene structure from a single RGB camera without the need for specialized depth sensors [25]. However, MDE models often struggle with ambiguities arising from textureless surfaces, specular reflections, and occlusions, all of which are especially prevalent in robotic surgical settings [4].
One avenue to address these limitations is to combine vision-based predictions with proprioceptive measurements [7][13]. We define robotic proprioception as a robot’s ability to sense contact with external objects in addition to an awareness of its own position in space through tool-tip kinematics. Utilizing proprioception, we can provide sparse ground-truth depth measurements at contact points between the robot and the environment. While these measurements can provide accurate estimates of depth at specified locations, their acquisition has a cost as each proprioceptive measurement requires physical interaction with the environment that will slow down task execution. Thus, we seek to minimize the number of required measurements while maximizing the reconstruction accuracy of the scene.
@@ II-A Monocular Depth Estimation
Monocular depth estimation uses an image from a single camera viewpoint to predict the distance from the camera to each point in the 3D scene. State-of-the-art MDE methods produce depth maps with outstanding accuracy for a wide variety of everyday indoor and outdoor scenes [37], [11]. However, endoscopic videos of surgical scenes are out of distribution from the large-scale datasets these foundation models are trained on, resulting in suboptimal performance [31].
Efforts to improve MDE in endoscopy include adapting or fine-tuning foundation models with simulation data [31], [32], [18], using a style transfer network to make simulation renderings more photo-realistic [10], or utilizing point cloud data from structure from motion as sparse-ground-truth depth cues [20]. MDE models focused on a specific domain frequently use an adaptation of the U-Net architecture [29], [10]. In this work, we draw from a similar toolbox of approaches while adding the ability to adaptively improve depth prediction through uncertainty-guided active sensing. As in prior works, we utilize a style-transfer network to make simulation RGB renderings more realistic. We then use this data to fine-tune a large foundation model, Depth Anything V2 [37], which we subsequently use to generate pseudo-ground-truth data for training a model with better real-world performance. We introduce a novel source of cues for endoscopic MDE, robotic proprioception at specific locations within the scene which will maximally resolve the uncertainty of our prediction.
@@ II-B Uncertainty Quantification
As deep neural networks become more complex and less interpretable, uncertainty quantification has become a critical component of any such system, especially for safety-critical tasks such as robotic surgery. Many methods exist for quantifying predictive uncertainty in deep learning including deep ensembles [17], Monte Carlo dropout [24], [2], [21], and test-time augmentation [34]. A comparison of these methods shows that deep ensembles consistently provide the most accurate quantification of uncertainty under distribution shift in the data [26]. An ensemble is created by training multiple models with different random weight initializations such that they each converge to a different local minima in the loss space. At inference time, we can use the variance of the ensemble members’ predictions as a proxy for uncertainty  [8][28][15]. Many works have quantified the uncertainty of MDE predictions [35][22][16], including some focused specifically on endoscopic images [3]. However, no past works have explored the active refinement of MDE predictions using robotic touch sensing. In this work, we leverage uncertainty quantification techniques for ensembles to develop an uncertainty-guided active touch sensing framework for monocular depth estimation.
@@ II-C Active Touch Sensing
Active touch sensing has been widely explored as a means of acquiring task-relevant information under sensing constraints. Early approaches focused on reconstructing object geometry through sparse and targeted contact interactions For instance, tactile fingertips have been used to actively trace object contours for shape recognition [1], while force-based probing has enabled reconstruction of unknown objects [12]. More recent work has emphasized uncertainty-driven exploration, where sensing actions are guided to reduce ambiguity [23]. Recent research has further advanced active touch sensing with machine learning and task-specific exploration. Learning-based approaches have been proposed to determine not only how to explore unknown objects, but also learning where to touch and how to recognize objects from tactile data [30, 36]. Together, these works demonstrate the value of active touch sensing to improve perception and task performance. In contrast, our work extends the concept of active touch sensing into the domain of proprioceptive depth acquisition, where sparse depth measurements are strategically chosen to reduce uncertainty in monocular depth estimation for deformable surgical environments.
@@ IV-A Monocular Depth Estimation
Our MDE is based on a U-Net architecture which takes as input the RGB image of a surgical scene as well as a sparse set of ground-truth depth measurements [29]. The sparse measurements are represented as a depth map where unobserved pixels are set to -1, enabling the network to condition its predictions on available propriocepted depths. The MDE then predicts a full depth map of the scene, leveraging both visual features in the RGB image as well as the sparse depth information.
@@ IV-C Active Depth Sensing
After computing the gradient of the total ensemble uncertainty with respect to the sparse depth measurements, $\nabla_{\mathrm{S}}U_{\mathrm{total}}$, we use this information to actively select which locations to propriocept. Rather than selecting points greedily, we adopt Stein Variational Gradient Descent (SVGD) to optimize a set of candidate locations [19]. Selecting points greedily could lead to a set of query points all near each other. SVGD strikes a balance between clustering points in highly informative regions while also encouraging diversity throughout the image. We first negate the gradient map so that the maximum values in the map correspond with locations with the highest likelihood of reducing uncertainty. The resulting map is then normalized to form a discrete probability distribution over all candidate pixel locations, indicating where measurements would be the most informative.
@@ V-A Datasets
Simulated Dataset: A simulated dataset is generated from CT scans of 13 real-world ex vivo porcine tracheal models with an intraluminal tumor surrogate (chicken tissue) affixed using cyanoacrylate adhesive. We convert the scans into 3D meshes in Unity. A simulated camera explores the interior of the trachea reconstructions, collecting pairs of RGB images and depth maps. The RGB images are then translated into the style of real phantom endoscopic scenes using a style-transfer network [27] to reduce the domain gap.
Real-world Dataset: Our real-world dataset is collected from 12 other ex vivo porcine tracheal models with tumor surrogates prepared as above. We collect exploratory endoscopic videos of these models using the Virtuoso Endoscopy System (Virtuoso, Nashville, TN, USA)  [33] and then generate pseudo-ground-truth depth maps using Depth Anything V2’s indoor large model [37] fine-tuned on the simulated dataset. We create additional data points by linearly and exponentially stretching the values of the pseudo-ground-truth depth maps. This artificially exaggerates the ambiguity of perceived scale, thereby forcing the model to rely heavily on the sparse ground truth.
@@ V-C Experimental Setup
Whereas in the simulated setting, a dense ground-truth depth map was readily available, this information does not exist in the physical system. Consequently, depth estimates are obtained through a touch-sensing tool integrated at the distal end of the concentric tube arm. The touch-sensing tool, first introduced in [13], employs a capacitance circuit coupled to an electrosurgery probe running along the arm. Tissue contact produces a sharp change in capacitance, which can be measured in real time, thereby enabling the arm to detect contact events. By using this tool to physically ‘probe’ image-space points, we are able to recover their corresponding depths via robot kinematics.
@@ V-E Evaluation Metrics
To evaluate the reconstruction quality of each method, we report standard monocular depth estimation metrics widely used in prior work [6, 5, 9, 14]: